OpenStudy (anonymous):

an investment is worth $2162 in 2007.By 2011 it has grown to $3810. Let y be the value of the investment in the year x, where x=0 represents 2007.Write a linear equation that models the value of the investment in the year x.

OpenStudy (tkhunny):

It's linear. That's code for changing the same amount at each stage. 2007 2162 2008 2162 + One Chunk 2009 2162 + two chuncks 2010 2162 + three chunks 2011 2162 + four chunks = 3810 2011 is the key 2162 + 4(Chunk) = 3810 Solving 4(Chunk) = 3810 - 2162 = 1648 Chunk = 1648/4 = 412 Can you now write the linear formula suggested?

OpenStudy (anonymous):

yes now i see thanks